HUBBLE SPACE TELESCOPE DAILY REPORT #3055 PERIOD COVERED: 0000Z (UTC) 02/12/02 - 0000Z (UTC) 02/13/02 Daily Status Report as of 044/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.2 Completed Two Sets of WF/PC-2 9060 (Photometry of a Statistically Significant Sample of Kuiper Belt Objects) The WF/PC-2 was used to propel the physical study of KBOs forward by performing accurate photometry at V, R, and I on a sample of up to 150 KBOs. The sample is made up of objects that will be observed at thermal infrared wavelengths by SIRTF and will be used with those data to derive the first accurate diameters and albedos for a large sample of KBOs. The observations completed nominally. 1.3 Completed Four Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.4 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.5 Completed STIS/CCD 9066 (Closing in on the Hydrogen Reionization Edge of the Universe) The Space Telescope Imaging Spectrograph (CCD) was used in parallel constrain the Hydrogen reionization edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted redshifts. The proposal completed uneventfully. 1.6 Completed WF/PC-2 9236 (Development And Evolution Of YSO Outflows And Jets) The WF/PC-2 was used to observe XZ Tauri, HL Tauri, and HH 30, three young stars which can be observed simultaneously in a single pointing. Broad-band HST imaging has revealed that the XZ Tauri binary is the source of an elongated bubble of emission nebulosity nearly 5" long. In new images of these objects obtained in March 1998, it was discovered that the XZ Tau bubble had expanded significantly and had become markedly limb-brightened since 1995. This behavior suggests that we are witnessing, for the first time, the earliest stages of development of a Herbig-Haro bowshock, perhaps the initial formation of its post-shock cooling zone. It is proposed to test this hypothesis with emission line imaging and temporal monitoring of the bubble's size and brightness. Simultaneously with the XZ Tau study, we will observe the jets and reflection nebulosities of HH 30 and HL Tauri. Monitoring of HH 30 is particularly important, as our 1998 imaging also revealed that a very large reflected light asymmetry has developed in this prototype young stellar object accretion disk system. There was no anomalous activity. 1.7 Completed STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.8 Completed STIS/CCD/MA1 9054 (Star Formation Triggers and Chemical Reprocessing in I Zw 18) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe I Zw 18, a blue compact galaxy {BCD} with the smallest known abundance of heavy elements yet observed. No anomalies were reported. 1.9 Completed Two Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.10 Completed Two Sets of STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.11 Completed Three Sets of STIS/MA1/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.12 Completed STIS/CCD/MA1/MA2 9036 (An Ultraviolet Spectroscopic Survey of Star-Forming Galaxies in the Local Universe) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to perform a comprehensive STIS ultraviolet spectroscopic survey of star-forming galaxies in the local universe. The sample covers a broad range of morphologies, chemical composition, and luminosity. The observations will provide spectral coverage between 1200 and 3100 Angstrom, at a resolution of 100 to 200 kms and S/N of about 30. The data set will allow the proposers to document and quantify the effects of massive stars on the interstellar medium and to infer implications for the evolution of the host galaxies. The proposal completed with no anomalous activity. 1.13 Completed STIS/CCD/MA2 9161 (The Ultraviolet Properties of Obscured QSOs) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to observe a subset of The Two Micron All-Sky Survey {2MASS} -discovered QSOs to: 1} determine if UV absorption- line features are present as seen in several other highly polarized AGN populations; 2} test relationships between X-ray and UV/optical attenuation, AGN spectral type, optical polarization, and detected UV absorption-line features; and 3} characterize the spectral energy distributions and UV properties of obscured QSOs. There were no problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 14 Successful: 14 2.3 Operations Notes: Joint Integrated Simulation (JIS) #5 (EVA-4), the last JIS in preparation for a planned February 28 Servicing Mission 3B launch, was conducted yesterday from 09:30 a.m. until 05:30 p.m. Using ROP SR-1A, the SSR-1 EDAC error counter was cleared twice. SSA transmitter #1 was turned on for the interval 043/235500Z to 044/000841Z per ROP IC-2. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW